The following budget data are for a country having both a central government and provincial governments:

Central purchases of goods and services 200

Provincial purchases of goods and services 150

Central transfer payments 100

Provincial transfer payments 50

Grants in aid (central to provincial) 100

Central tax receipts 450

Provincial tax receipts 100

Interest received from private sector by central government 10

Interest received from private sector by provincial governments 10

Total central government debt 1000

Total provincial government debt 0

Central government debt held by provincial governments 200

Nominal interest rate 10%

Calculate the overall and primary deficits for the central government, the provincial governments, and the combined governments.


1
Expert's answer
2021-11-26T12:15:36-0500

Ans - The primary deficit for the combined governments is -$50

Explanation:

Primary deficit for central govt = Govt expenses + Transfer − Taxes

Primary deficit for central govt = Govt expenses + Transfer - Taxes

= Central purchases of goods and service + Central transfer payments +

 Grants in aid​ (central to​ provincial) − Central tax receipts

= Central purchases of goods and service + Central transfer payments + Grants in aid​ (central to​ provincial) - Central tax receipts

= 200 + (100 + 100) - 450

= -50

Primary deficit for provincial = Govt expenses + Transfer − Taxes

Primary deficit for provincial = Govt expenses + Transfer - Taxes

= 150 + 50 -100 

= 100

Primary deficit for combined govt = -50+ 100 = -$50

   =  Central purchases of goods and service + Central transfer payments +

 Grants in aid​ (central to​ provincial) + Interest receive − interest paid 

− Central tax receipts

   =  Central purchases of goods and service + Central transfer payments + Grants in aid​ (central to​ provincial) + Interest receive - interest paid - Central tax receipts

= 200 + (100 + 100)+ 10 - (1000*10%) - 450

= -140

Overall deficit for provincial  = Govt expense + Transfers + Interest −

 Taxes

Overall deficit for provincial  = Govt expense + Transfers + Interest - Taxes

150 + 100 -100 + 10 - ( 450*10%) - 100

= 15

Overall deficit of combined govt = -140 + 15 = -125